### v1

# Towards Verifying Exact Conditions of Density Functional Theory Approximations

Abstract
Density Functional Theory (DFT) is used extensively in the computation of electronic properties of matter, with applications in solid state physics, computational chemistry, and materials science. Approximating the exchange-correlation (XC) functional is the key to the Kohn-Sham DFT approach, the basis of most DFT calculations. The choice of this density functional approximation (DFA) depends crucially on the particular system under study, which has resulted in the development of hundreds of DFAs. Though the exact density functional is not known, researchers have discovered analytical properties of this exact functional. Furthermore, these exact conditions are used when designing DFAs.
This paper presents XCVerifier, the first approach for verifying whether a DFA implementation satisfies the DFT exact conditions. XCVerifier was evaluated on five DFAs from the popular Libxc library and seven exact conditions used in recent work by Pederson and Burke. XCVerifier was able to verify or find violations for a majority of the DFA/condition pairs, demonstrating the feasibility of using formal methods to verify DFA implementations. However, it timed out on all conditions of the recent SCAN functional, revealing potential directions for future work.
Index Terms:
Density functional theory, Formal methods.
I Introduction
Density functional theory (DFT) [1, 2, 3] is a widely used approximation in the computation of electronic properties of matter. DFT calculations have been used to predict diverse properties–from atomic binding energies and chemical reactivities to electronic conductivities and magnetic phenomena [4]. Consequently, DFT finds applications in a number of scientific and engineering fields, including solid state physics, computational chemistry, and materials science, and has been implemented in widely-used scientific software such as ABINIT [5], cp2k [6], ERKALE [7], Psi4 [8], Octopus [9], Qbox [10], and Quantum ESPRESSO [11].
Originating in the foundational work of Hohenberg and Kohn in 1964 [12], DFT has provided a practical way to reduce the complexity of conventional electronic structure methods by identifying the electron density function $n(r)$—a real function in three-dimensional space—as the fundamental quantity from which all other properties of a physical system can be derived. The Kohn-Sham (KS) [13] approach is currently the basis of most DFT calculations. The KS DFT approach states that only the exchange-correlation (XC) energy portion of a functional needs to be approximated. However, the exact expression for this XC energy functional $E_{\rm xc}[n]$, which describes the complex electron-electron interactions within systems, is not known and is incredibly difficult to approximate [14].
Researchers have developed hundreds of approximations to the XC energy functional, which have the form:
[…]
where the term $\tilde{\epsilon}_{\text{xc}}$ is the density functional approximation (DFA). The choice of DFA in a particular application depends crucially on the known or expected physical properties of the system under study [15, 16, 17].
This proliferation of DFAs is reflected in the fact that the popular Libxc software library [18], which provides numerical implementations of DFAs, currently includes over 500 functionals [19]. The DFAs are also of varying degrees of complexity [20]: (i) local density approximations (LDAs) depend only on the electron density $n$, (ii) generalized-gradient approximations (GGAs) [21] depend on the electron density $n$ and its gradient $\nabla n$, and (iii) meta-GGAs have a further dependence on the Laplacian $\nabla^{2}n$ and the local kinetic energy density $\tau$.
The DFA $\tilde{\epsilon}_{\text{xc}}$ in Equation 1 is a mathematical function with a known, albeit complicated, analytical form. For instance, the following equation shows just the exchange (X) part $\tilde{\epsilon}_{\text{x}}^{\,\text{PBE}}$ of the Perdew-Burke-Ernzerhof (PBE) functional [22], a commonly used GGA functional [23]:
$\tilde{\epsilon}_{\text{x}}^{\,\text{PBE}}(\rho,\sigma)=\frac{2.884\rho^{1/3}\left(-28.944\pi^{4/3}\rho^{8/3}-0.174\pi^{2}\sigma\right)}{\pi^{1/3}\cdot\left(77.184\pi^{4/3}\rho^{8/3}+0.257\pi^{2}\sigma\right).}$
However, the correlation part of PBE is significantly more complex with over 300 operations in the Libxc implementation. The SCAN meta-GGA functional [24] is even more complex with over 1000 operations in the exchange and correlation parts combined, including transcendental functions such as $\exp$ and $\log$.
Creating a new DFA is an art that has only been mastered by few researchers as of today. The functional forms used in the definition of a DFA vary considerably. DFA designs fall into two categories: empirical and non-empirical. Empirically-designed DFAs (e.g., LYP [25]) are tailored for molecular chemistry applications and perform well on molecular benchmarks [26, 27]. Non-empirically designed DFAs are constructed to satisfy some exact conditions, which are known analytical properties of the exact functional (Section II). For example, the correlation energy $E_{C}$ non-positivity condition states that the correlation energy cannot be positive, i.e., $E_{\rm c}[n]\leq 0$ [28]. Furthermore, so-called norms are imposed on DFAs by requiring that they reproduce correctly some known physical systems, e.g. a hydrogen or a helium atom for which exact results are available. The SCAN functional is built to satisfy as many as 17 constraints and norms [24].
Recently, Pederson and Burke (PB) [28] checked whether the Libxc implementations of various DFAs satisfy DFT exact conditions. In particular, they checked seven exact conditions by considering their corresponding local condition (Section II). These local conditions were assessed for a given DFA by performing a grid search over the input variables, and checking whether each input-output pair satisfies the local condition. The PB approach is the state of the art in the DFT community, and was the first to perform a large scale study of the role of exact conditions in density functional development. For instance, they found that many empirical DFAs satisfy these exact conditions in certain regions even though they were designed without explicit adherence to these exact conditions. Many of the conditions require gradient or limit calculations, which need to be numerically approximated.
This paper addresses the problem of automatically verifying whether the implementation of a DFA satisfies the exact conditions of the density functional. It is the first to apply formal-methods techniques to density functional theory. The aim is to provide formal guarantees related to the correctness of existing DFAs; viz., to formally verify if the DFA satisfies the exact conditions, and to determine the input regions in which it does not. The prior PB approach might miss violations of the property (because of the use of sampling) or report violations are not true (because many local conditions require gradient or limit calculations that were numerically approximated).
As a step towards solving this problem, we designed and implemented XCVerifier, a tool that automatically converts a Libxc functional (implemented in Maple[29]) to Python, uses dReal [30] and SymPy [31] to compute any required derivatives and limits in the local condition, and then uses the dReal solver to verify that the DFA satisfies the condition in the relevant input region (Section III). We also implemented a domain-splitting technique to improve the performance of XCVerifier (Algorithm 1).
We evaluated XCVerifier by verifying seven exact conditions (from Pederson and Burke) for five popular DFAs: PBE [22], SCAN [24], LYP [25], AM05 [32], and VWN RPA [33], which cover the different types of DFAs (LDA, GGA, and meta-GGA), as well as different design categories (empirical and non-empirical). Some conditions do not apply to certain DFAs, which left us with 29 DFA/condition pairs. As shown in Section IV-B, XCVerifier was successfully able to verify or find a counterexample for 11 pairs, and it was able to verify all but a small portion of the input domain for an additional 7 pairs (Table I). These results demonstrate the feasibility of using formal-methods techniques to verify DFA implementations. However, XCVerifier timed out for 11 pairs: it was unable to verify the $U_{\text{c}}$ monotonicity for all but the simplest of the DFAs we considered (VWN RPA), and was unable to verify any of the properties for the SCAN DFA. This motivates further research on improving formal-methods techniques in the context of DFT, an important scientific computation discipline.
We also compared the results from the PB approach with XCVerifier (Section IV-C). We found that the results of the PB approach were inconsistent with those of XCVerifier for 5 DFA/condition pairs, all of which were related to the two $T_{\text{c}}$ upper bound conditions (Table II). For instance, the PB approach found no violations to the $T_{\text{c}}$ upper bound condition for the PBE functional (Figure 2(b)), while XCVerifier was able to find a number of violations (Figure 2(e)). The $T_{\text{c}}$ upper bound condition entails computing derivatives and limits for which the PB approach uses numerical approximations. In contrast, XCVerifier computes these symbolically. This demonstrates one advantage of our approach to verifying DFT exact conditions.
The contributions of the paper are as follows:
  • A tool, XCVerifier, for automatically verifying exact conditions for density functional approximations (Section III).
  • An evaluation of XCVerifier using five DFAs and seven exact conditions along with a comparison with the state-of-the-art grid-search approach (Section IV).
Our preliminary results demonstrate the feasibility of using formal methods to improve the correctness of DFT implementations, and reveal potential avenues for future work (Section VI).
II Exact Conditions in DFT
This section lists the exact conditions of the density functional $E_{\rm xc}$ considered in this paper. Our description closely follows that in Pederson and Burke [28]. Each exact condition has a corresponding local condition such that if the DFA $\tilde{\epsilon}_{\text{xc}}$ satisfies the local condition, then the functional $\widetilde{E}_{\rm xc}$ satisfies the (global) exact condition.
While the exact conditions take as input the global energy $E_{\rm xc}[n(r)]$ of the DFA, the local conditions take the exchange (correlation) enhancement factor $\widetilde{F}_{\text{xc}}$, which is a function of the local value ${\epsilon}_{\text{xc}}\big{(}n(r)\big{)}$ of the exchange (correlation) energy of the DFA. For GGA functionals, the inputs $n(r)$ and $\nabla n(r)$ are usually expressed in terms of the Wigner-Seitz radius $r_{s}=(r\pi n/3)^{-1/3}$ and $s=|\nabla n|/(2(3\pi^{2})^{1/3}n^{4/3})$. Given the expressions for the exchange and correlation energies $\tilde{\epsilon}_{\text{x}}$ and $\tilde{\epsilon}_{\text{c}}$ of the DFA in terms of $s$ and $r_{s}$, we can compute $\widetilde{F}_{\text{x}}$ and $\widetilde{F}_{\text{c}}$ of $\widetilde{F}_{\text{xc}}$:
$\widetilde{F}_{\text{xc}}[n(r)]=\widetilde{F}_{\text{x}}+\widetilde{F}_{\text{c}}=\frac{\tilde{\epsilon}_{\text{xc}}[n(r)]}{{\epsilon}_{\text{x}}^{\,\text{unif}}[n(r)]}.$ (2)
We now list the DFT exact conditions along with their corresponding local condition:
  • The correlation energy $E_{\rm c}$ non-positivity condition [28] is defined as
    $E_{\rm c}[n]\leq 0.$ (3)
    The corresponding local condition is
    ${\epsilon}_{\text{c}}(n(r))\leq 0.$ (4)
  • The $E_{\rm c}$ scaling inequality [34, 28],
    $(\gamma-1)E_{\rm c}[n_{\gamma}]\geq\gamma(\gamma-1)E_{\rm c}[n],$ (5)
    has the local condition
    $\frac{\partial\widetilde{F}_{\text{c}}}{\partial r_{s}}\geq 0.$ (6)
  • The $U_{\text{c}}(\lambda)$ monotonicity condition [35, 28] is
    $\frac{dU_{\text{c}}(\lambda)}{d\lambda}\leq 0,$ (7)
    where $U_{\text{c}}(\lambda)=d(\lambda^{2}E_{\rm c}[n_{1/\lambda}])/d\lambda$ represents the correlation energy adiabatic connection curves. The corresponding local condition is
    $\frac{\partial^{2}\widetilde{F}_{\text{c}}}{\partial r_{s}^{2}}\geq\frac{-2}{r_{s}}\cdot\frac{\partial\widetilde{F}_{\text{c}}}{\partial r_{s}}.$ (8)
  • Lieb-Oxford bound [36, 28] is
    $U_{\text{xc}}\geq C_{\text{LO}}\int d^{3}r\,n(r)\,{\epsilon}_{\text{x}}^{\,\text{unif}}\big{(}n(r)\big{)}.$ (9)
    where $U_{\text{xc}}[n]=E_{\rm xc}[n]-T_{\text{c}}[n]$ is the potential correlation energy and $C_{\text{LO}}=2.27$ is the Lieb-Oxford constant, following [28]. The corresponding local condition is
    $\widetilde{F}_{\text{xc}}+r_{s}\frac{\partial\widetilde{F}_{\text{c}}}{\partial r_{s}}\leq C_{\text{LO}}.$ (10)
  • The Lieb-Oxford extension to $E_{\rm xc}$ [36, 28] is a generalization of the Lieb-Oxford bound with $E_{\rm xc}$ instead of $U_{XC}$:
    $E_{\rm xc}\geq C_{\text{LO}}\int d^{3}r\,n(r)\,{\epsilon}_{\text{x}}^{\,\text{unif}}\big{(}n(r)\big{)}.$ (11)
    The corresponding local condition is
    $\widetilde{F}_{\text{xc}}\leq C_{\text{LO}}$ (12)
  • The $T_{\text{c}}$ upper bound condition [35, 28] is
    $T_{\text{c}}[n_{\gamma}]\leq-\gamma\left(\frac{\partial E_{\rm c}[n_{\gamma}]}{\partial\gamma}\right)+E_{\rm c}[n_{\gamma}].$ (13)
    The corresponding local condition is
    $\frac{\partial\widetilde{F}_{\text{c}}}{\partial r_{s}}\leq\frac{\widetilde{F}_{\text{c}}(\infty)-\widetilde{F}_{\text{c}}}{r_{s}}.$ (14)
    where $\widetilde{F}_{\text{c}}(\infty)$ is the the limit of $\widetilde{F}_{\text{c}}$ as $r_{s}\to\infty$.
  • The conjectured $T_{\text{c}}$ upper bound [37, 38, 39, 28] is
    $T_{\text{c}}[n]\leq-E_{\rm c}[n]$ (15)
    with local condition
    $\frac{\partial\widetilde{F}_{\text{c}}}{\partial r_{s}}\leq\frac{\widetilde{F}_{\text{c}}}{r_{s}}.$ (16)
III XCVerifier
This section describes the design the XCVerifier tool that verifies whether an DFA implemented in the Libxc library [18] satisfies the DFT exact conditions. XCVerifier consists of two main parts: XCEncoder, which encodes the given exact condition for a given Libxc functional into a formula $\psi$, and Verifier, which verifies whether this encoded formula $\psi$ is always true (valid) in the given input domain. Because DFAs involve non-linear arithmetic as well as transcendental functions, we chose the dReal solver [30] as the core solver in the Verifier. As a consequence, the formula $\psi$ generated by XCEncoder is a dReal formula.
III-A XCEncoder
Given a Libxc functional, XCEncoder first translates the Maple code for $\tilde{\epsilon}_{\text{xc}}$ to Python using CodeGeneration package of Maple. To test that this translation is correct, we uniformly sample points from the input region and assert that the outputs of the translated Python implementation and the original Libxc implementation are close by a factor of at least $10^{-13}$. Of the DFAs we use in our experiments, all points for PBE, AM05, and VWN RPA are within $10^{-14}$ of the Libxc implementation. LYP is within $10^{-13}$ and SCAN is within $10^{-14}$ for $98\%$ of the points sampled.
The next step in the XCEncoder is to convert the Python code into a dReal expression. We implemented a symbolic execution engine for (a subset of) Python that generates this dReal expression. Note that though DFA implementations do not contain loops, arrays, etc., they do contain (non-recursive) function calls and if-then-else statements.
XCEncoder needs to return the dReal formula $\psi$ that encodes the given exact condition for the particular functional. Encoding the local condition corresponding to the $E_{\rm c}$ non-positivity condition is straightforward: XCEncoder uses the dReal expression for $\tilde{\epsilon}_{\text{c}}$ to directly construct the dReal formula $\tilde{\epsilon}_{\text{c}}\leq 0$.
However, the local conditions corresponding to exact conditions such as $E_{\rm c}$ scaling, Lieb-Oxford, $U_{\text{c}}$ monotonicity, and $T_{\text{c}}$ upper bound require computation of one or more derivatives, In such cases, XCEncoder uses the symbolic differentiation built into dReal to compute the derivatives. Furthermore, the local condition corresponding the $T_{\text{c}}$ upper bound condition requires computing a limit. In such cases, XCEncoder uses SymPy [31] to compute the limit, then translates this limit back to a dReal expression. After any such derivatives and limits are symbolically computed, XCEncoder constructs the required formula $\psi$ corresponding to the specific local condition.
III-B Verifier
Algorithm 1 Verifier(input domain $D$, formula $\psi$)
1:  if $D<t$ then
2:     return
3:  result, x $\leftarrow$ dReal($\varphi_{D}\wedge\lnot\psi$)
4:  if result $=$ UNSAT then
5:     print  “Verified condition over domain $D$
6:     return
7:  if result $=$ SAT then
8:     if valid(x) then
9:        print  “Found counterexample x”
10:     else
11:        print  “Verification inconclusive over domain $D$
12:  else
13:     print  “Verification timed out over domain $D$
14:  for all $D^{\prime}$ in split($D$do
15:     Verifier($D^{\prime}$, $\psi$)
16:  return
The Verifier takes as input the formula $\psi$ encoding the exact condition for the DFA and a domain for the inputs to the DFA. We use the same input bounds as used in Pederson and Burke [40]; for example, for GGA functionals, the domain for $r_{s}$ is the interval $[0.0001,5]$, and that for $s$ is $[10^{-12},5]$ (to avoid division-by-zero errors). Consequently, for a GGA functional, the Verifier is trying to prove the validity of the following formula:
$\forall r_{s},s\,.\ (r_{s}\in[0.0001,5]\wedge s\in[10^{-12},5])\implies\psi.$ (17)
The validity of Equation 17 is equivalent to proving that the following formula is unsatisfiable:
$r_{s}\in[0.0001,5]\wedge s\in[10^{-12},5]\wedge\lnot\psi.$ (18)
[…]
Our preliminary results showed that dReal would timeout (in 2 hours) when trying to solve the formula in Equation 18. To improve performance of Verifier, we implemented a domain-splitting technique that partitions the input domain and use dReal to solve the formula $\lnot\psi$ on each subdomain separately. This simple strategy greatly improved the performance of Verifier. Furthermore, one of the goals of XCVerifier is to determine the input regions where the DFA does not satisfy the exact condition. Thus, Verifier also performs the domain-split when dReal returns a valid input violating the given condition. This allows Verifier to isolate the regions where the DFA does not satisfy the local condition.
Algorithm 1 presents the pseudo-code for Verifier, and Figure 1(e) shows a graphical representation of output of Verifier when verifying the $E_{\rm c}$ scaling inequality condition for the LYP functional. The dReal solver is called on Line 3 to find a satisfying assignment to the formula $\varphi_{D}\wedge\psi$, where $\varphi_{D}$ is the formula encoding the domain constraints on the inputs to the DFA. If the result is UNSAT, then Verifier returns after recording that the condition was verified over domain $D$ (Line 6), indicated using   in Figure 1(e). If the result is SAT and the returned model is a valid model for $\lnot\psi$, then Verifier records that a valid counterexample was found (Line 9), else the result is inconclusive (Line 11), indicated using   and  , respectively, in Figure 1(e). It is also possible that dReal is unable to determine satisfiability within the two hour time limit. In this case, Verifier records that the verification timed out for domain $D$ (Line 13), indicated using   in Figure 1(e).
Line 14 is executed if the result is SAT or the solver timed out. Verifier calls split($D$), which partitions each input dimension of $D$ into two equal parts. The Verifier is recursively called on each subdomain $D^{\prime}$ on Line 15. We set a lower limit on the size of input domain as the base case for the recursion. On Line 2, Verifier returns if the given input domain is too small as determined by the threshold $t$. For our experiments, we set the limit for recursing after a counterexample to $0.2$ for the widths of the $r_{s}$ and $s$ ranges. For an inconclusive or timeout result, we set it to $0.625$.
TABLE I: Verifying local conditions for DFT exact conditions for DFAs using XCVerifier.
: XCVerifier verified that the DFA satisfies the condition on the entire input domain;
[…]
: XCVerifier found a counterexample showing that the DFA does not satisfy the condition.
Local condition PBE LYP AM05 SCAN VWN RPA
[…]
IV Experimental Results
This section evaluates the performance of XCVerifier for verifying exact conditions for DFAs, and compares the results to the prior PB approach [28].
Our experiments were designed to answer the following research questions:
  1. RQ1
    Is XCVerifier able to verify or find counterexamples for DFAs (Section IV-B )?
  2. RQ2
    How does compare XCVerifier to the PB approach (Section IV-C)?
IV-A Experimental Setup
We use the following five DFAs: PBE [22], a popular non-empirical GGA DFA; SCAN [24], a fully constrained non-empirical meta-GGA DFA satisfying all known properties of DFAs; LYP [25], and empirical DFA that is a key component of several commonly-used DFAs; AM05 [32], which shows efficient and superior performance on solids [41]; and VWN RPA [33], an LDA functional.
For each of these DFAs, we consider each of the applicable conditions from Section II. Note that the Lieb-Oxford conditions only apply to functionals with both an exchange and correlation component available (e.g. PBE, SCAN).
The PB approach is the state of the art for assessing condition satisfaction in DFT. For a given DFA and condition, the PB approach draws $10^{5}$ uniform samples each for $r_{s}$ and $s$, which are then meshed into a grid. PB then calls the Libxc implementation of the DFA for each of the points in the grid. This grid is used to numerically compute the limits and gradients necessary for the conditions using the NumPy package in Python [42]. Then the condition is checked at each point in the grid. The condition is assumed to be satisfied for the DFA if all the points in the grid pass the condition.
IV-B Verifying Local Conditions using XCVerifier
Table I summarizes the result of using XCVerifier to verify the local condition corresponding to each DFA exact condition in Section II for the five DFAs. Visualization of the results for PBE and LYP are shown in Figure 2 and Figure 1.
From Table I, we can see that XCVerifier rarely determines that a condition is fully satisfied () for a DFA: it determines that AM05 and VWN RPA satisfy the $E_{\rm c}$ non-positivity condition, and that VWN RPA satisfies the $E_{\rm c}$ scaling inequality condition.
For some of the DFAs, XCVerifier finds counterexamples to the conditions (). It finds counterexamples to the $T_{\text{c}}$ upper bound condition for all the DFAs except for SCAN, where it times out. It also finds counterexamples for the conjectured $T_{\text{c}}$ upper bound condition for PBE and LYP, and for all conditions but $U_{\text{c}}$ monotonicity for LYP.
[…]
Interestingly, XCVerifier also times out completely for every single condition applied to SCAN, which was designed to satisfy all known properties of DFAs.
PBE  For the PBE functional, XCVerifier finds a counterexample region at $r_{s}<3.125$ and $s>1.25$ for the conjectured $T_{\text{c}}$ upper bound (Figure 2(f)). It is able to verify the entire input domain for the Lieb-Oxford bounds except for at $r_{s}<0.625$, where it times out. This is also true for the $E_{\rm c}$ non-positivity condition, which also results in some inconclusive and timeout regions. For the $E_{\rm c}$ scaling inequality, XCVerifier times out in the entire domain except for where $r_{s}>0.625$ and $s>3.75$. For the $T_{\text{c}}$ upper bound condition, it finds counterexamples in $r_{s}>0.6250875$ and times out in the rest of the region.
LYP  For the $E_{\rm c}$ non-positivity condition, XCVerifier returns $s>1.25$ as the counterexample region (Figure 1(d)). For the $E_{\rm c}$ scaling inequality, it returns $r_{s}<2.19$ and $s>0.94$ (Figure 1(e)). It times out for the entire input domain for the $U_{\text{c}}$ monotonicity condition and for the $T_{\text{c}}$ upper bound condition, it verifies the part of the region where $1.8750.3125<r_{s}<$ 1.875 and $s>1.25$ and finds counterexamples in the rest of the region. For the conjectured $T_{\text{c}}$ upper bound condition, XCVerifier returns $r_{s}>0.3125$ and $s>0.9375$ as the failing region (Figure 1(f)). There are also some inconclusive and timeout regions for each of the conditions, usually bordering the counterexample regions.
AM05XCVerifier verifies that AM05 satisfies $E_{\rm c}$ non-positivity for the entire region and times out completely for $U_{\text{c}}$ monotonicity. For the $E_{\rm c}$ scaling inequality, it verifies the entire input region except for $r_{s}<0.625$, where it times out. It verifies the same region for the $T_{\text{c}}$ upper bound conjectured condition and returns inconclusive for $r_{s}<0.625$. It finds no counterexamples for any of the conditions.
SCAN  For the SCAN functional, XCVerifier times out after 2 hours for all of the conditions. To check if XCVerifier is able to verify a weaker condition for SCAN, we relax the the $E_{\rm c}$ negativity condition by adding a tolerance of $0.1$ such that the verifier needs to find a model for $E_{\rm c}<0.1$ rather than $E_{\rm c}<0$. However, XCVerifier times out completely for the weakened condition as well.
VWN RPAXCVerifier verifies that VWN RPA satisfies the $E_{\rm c}$ non-positivity and $E_{\rm c}$ scaling inequality for the entire input domain. For the $U_{\text{c}}$ monotonicity condition, it verifies the entire region except for where $r_{s}<0.001$, where it times out. For the $T_{\text{c}}$ upper bound, it finds counterexamples in almost the entire region except for some small subregions in the middle of the input region where it times out. For the conjectured $T_{\text{c}}$ upper bound, XCVerifier is inconclusive for $r_{s}<0.3125$. The remaining region is verified as satisfying the condition.
Summary for RQ1:  In our evaluation of the seven exact conditions for the five DFAs, XCVerifier was able to verify or find counterexamples for 11 condition-DFA pairs and partially verify 7, as shown in Table I. This demonstrates the feasibility of using formal-methods techniques like XCVerifier in verifying DFA implementations. However, XCVerifier also timed out for 11 of the 29 applicable pairs, particularly for the $U_{\text{c}}$ monotonicity condition and for the SCAN functional. Thus, there is room for improvement in formal methods techniques for DFT.
IV-C Comparing XCVerifier to PB
TABLE II: Comparison between results for XCVerifier and PB approach.
[…]
Local condition PBE LYP AM05 SCAN VWN RPA
[…]
(a) $E_{\rm c}$ non-positivity w. PB
(b) $E_{\rm c}$ scaling inequality w. PB
(c) Conj. $T_{\text{c}}$ upper bound w. PB
(d) $E_{\rm c}$ non-positivity w. XCVerifier
(e) $E_{\rm c}$ scaling inequality w. XCVerifier
(f) Conj. $T_{\text{c}}$ upper bound w. XCVerifier
Figure 1: Regions where the LYP functional satisfies or violates conditions according to PB (top) and XCVerifier (bottom).
For PB: ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\mathllap{\circ}$ is a counterexample to the condition, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bullet}\mathllap{\circ}$ is a point that satisfies the condition.
For XCVerifier:   is a region that contains a counterexample marked with ×,   is a region that is verified to satisfy the condition,   indicates a timeout, and   indicates an inconclusive result.
(a) $E_{\rm c}$ non-positivity w. PB
(b) $T_{\text{c}}$ upper bound w. PB
(c) Conj. $T_{\text{c}}$ upper bound w. PB
(d) $E_{\rm c}$ non-positivity w. XCVerifier
(e) $T_{\text{c}}$ upper bound w. XCVerifier
(f) Conj. $T_{\text{c}}$ upper bound w. XCVerifier
Figure 2: Regions where the PBE functional satisfies or violates conditions according to PB (top) and XCVerifier (bottom).
For PB: ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\mathllap{\circ}$ is a counterexample to the condition, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bullet}\mathllap{\circ}$ is a point that satisfies the condition.
For XCVerifier:   is a region that contains a counterexample marked with ×,   is a region that is verified to satisfy the condition,   indicates a timeout, and   indicates an inconclusive result.
This section compares the results for XCVerifier to the PB approach.  Table II shows the consistency between the results of PB and XCVerifier.
[…]
Inconsistent results, $\bigotimes$, mean that the counterexample regions returned by PB and XCVerifier differ significantly. This occurs for the $T_{\text{c}}$ upper bound condition for all of the functionals except for SCAN, where XCVerifier times out. In our experiments, the inconsistencies for this condition are dramatic: the regions found by PB and XCVerifier are sometimes completely opposed. For example, for the $T_{\text{c}}$ upper bound condition for PBE (Figure 2(e)), the PB finds no counterexamples, while XCVerifier finds counterexamples in the entire region except where it times out.
The $T_{\text{c}}$ upper bound condition is the only condition that requires a limit computation, which is done numerically in PB and symbolically in XCVerifier. This may explain the discrepancy. There is also a less dramatic inconsistency for the conjectured $T_{\text{c}}$ upper bound condition for PBE Figure 2(f).
Several of the results are not inconsistent, marked with $\bigodot^{*}$, meaning that neither method finds counterexamples for the condition. This is the case for when PB finds no counterexamples and concludes that the condition is satisfied for the functional, and XCVerifier either verifies the entire region or partially times out and partially verifies the region. In general, this tends to looks like XCVerifier verifying the majority of the region with a smaller timed out or inconclusive region on the border of the input domain. This is the case for the $E_{\rm c}$ non-positivity condition (e.g. for PBE Figure 2(d)), the $E_{\rm c}$ scaling inequality condition, and the Lieb-Oxford conditions when XCVerifier does not time out completely.
[…]
Summary for RQ2:  Our comparison of XCVerifier to the PB approach (Table II) shows inconsistencies for 5 DFA/condition pairs, all of which were related to the two $T_{\text{c}}$ upper bound conditions. The $T_{\text{c}}$ upper bound condition requires computing limits of $\widetilde{F}_{\text{c}}$, which is a function of the potentially complicated correlation energy $\tilde{\epsilon}_{\text{xc}}$. XCVerifier computes this limit symbolically. In contrast, the PB approach uses numerical approximation, which may explain the discrepancy. This demonstrates one advantage of our approach to verifying DFT exact conditions.
V Related Work
To the best of our knowledge, ours is the first work that uses formal methods to verify correctness in the context of density functional theory, an important scientific computing application. Prior work has used a testing-based approach, which are more scalable but do not provide formal guarantees. As discussed earlier, Pederson and Burke [40] use a grid-search to evaluate whether the DFA satisfies the DFT exact conditions. Lehtola and Marques [43] demonstrate the numerical instability of many commonly-used DFAs by observing slow convergence of self-consistent field (SCF) calculation with typically used numerical quadrature schemes.
In general, correctness in scientific computing is recognized as a major challenge in HPC [44, 45], which needs formal methods that address the unique challenges in this domain. Progress to date includes the verification of mathematical properties in a conjugate gradient solver, a finite difference stencil, and a mesh quality metric [46], PDE solvers [47], specific properties of CG [48] and LU decomposition [49], and the floating-point equivalence of manually and automatically differentiated code [50]. The above are orthogonal to our goal of verifying exact conditions of density functional theory approximations.
Testing and analysis techniques for floating-point programs have been developed in the past decade. The first set of techniques are general approaches that aim to achieve high-coverage of numerical code  [51, 52], conduct differential testing of numerical libraries [53], and perform mutation testing of floating-point expressions given a real specification [54]. However, it has been shown that simply achieving high code coverage in numerical programs does not uncover numerical issues in most cases [52], and differential testing is not feasible for DFAs because most of them are unique.
A large body of work focuses on performing automated error analysis of floating-point programs [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]. While some of these approaches provide sound error bounds of floating-point programs, they suffer from important limitations with respect to program size and control structures supported. In the absence of techniques that can reason about floating-point error in non-trivial programs, a rich area of research in software testing has focused on how to efficiently generate inputs that maximize error in the output of a program  [67, 52, 68], which can shed light on the potentially worst error a floating-point program could incur in. Similarly, work has proposed techniques to generate inputs that trigger floating-point exceptions [69, 70]. Calculating error bounds of DFAs is orthogonal to our goal of verifying physical and numerical properties of their implementations.
Existing work has explored finding function input ranges, also referred to as regimes, with the purpose of improving the accuracy of floating-point expressions [71] or optimizing floating-point efficiency [72]. These approaches are based on either estimating error based on dynamic input sampling, or statically performing an error analysis, which have their own limitations, as described earlier.
VI Discussion
VI-A Improving scalability of the solver
XCVerifier was unable to verify any of the exact conditions of the SCAN functional, which has been designed to satisfy all known exact conditions. SCAN is significantly more complex than the other functionals we considered, and also involves the use of transcendental functions such as $\exp$ and $\log$. This causes dReal to timeout even when the input domain is reduced. It would be interesting to investigate approaches to improve the performance of the solver so that it can tackle the SCAN functional. Apart from its popularity, the SCAN functional will serve as a fascinating use case: there are a progression of DFAs—rSCAN, r++SCAN, r2SCAN, r4SCAN—proposed in the literature that were designed with different adherence to exact conditions to improve the numerical stability of the original SCAN functional [73, 74, 75].
VI-B Expanding to more DFA/condition pairs
Our evaluation demonstrated the robustness of the XCEncoder to handle a wide variety of DFAs and exact conditions. The ultimate goal of our research is to be able to analyze all the 500+ functionals in Libxc for all known DFT exact conditions. Future work will continue to expand our evaluation, and will aim to integrate our verification tool into the Libxc, e.g., as part of the continuous integration (CI) for Libxc.
VI-C Numerical issues with DFAs
Apart from verifying known exact conditions for DFA implementations, it would interesting to analyze numerical issues of the implementations with the goal of using formal methods to find and fix numerical issues in DFA implementations. This is a challenging problem involving reasoning about floating points and dealing with transcendental functions like $\sin$, $\log$ and $\exp$. The functional forms of DFAs themselves can also be a source of numerical issues. Some DFAs include different functions that apply to different input domains, and must ensure continuity when switching from one domain to another. Additionally, the parametrization of the DFA may cause issues. Even in the simple case of the Local Density Approximation (LDA), the Perdew-Zunger [76] parametrization of the results of Ceperley and Alder [77] includes potentially inaccurate numerical constants that lead to discontinuities of the exchange-correlation energy at a given matching point.
The functional form of a DFA may also make it sensitive to inaccuracies in its input data. While a given implementation of a DFA may yield correct answers for an exactly known (e.g. exponential) density, it may result in large numerical errors if the input density is noisy or if the density and its gradient are not numerically consistent. This is particularly problematic in regions of low density, e.g. far from a molecule placed in vacuum. Such large errors may lead to inaccurate energies or slow convergence in the solution of the Kohn-Sham equations. For example, the sensitivity of the SCAN functional requires the use of extremely fine grids to represent the electron density in order to avoid large numerical errors. This led some authors to modify the SCAN functional to avoid this numerical issue, resulting in a slightly different DFA [74]. In other cases, an analytical reformulation of a DFA is used to avoid numerical issues [78] without modifying it. These fixes are however ad hoc, and there is no known general recipe for avoiding the numerical issues of a DFA.
VII Conclusion
This paper presented XCVerifier, an approach for verifying whether a DFA implementation satisfies DFT exact conditions. XCVerifier automatically encodes the DFA implementation from Libxc and a given exact condition into a dReal formula, symbolically performing any required derivative and limit calculations. It uses the dReal solver to verify whether the condition is true or find a violation to it for a given input domain. XCVerifier also implements a domain-splitting technique to improve performance by reducing solver timeouts, and isolate the input regions where the condition is satisfied or violated.
We evaluated XCVerifier by verifying seven exact conditions (from Pederson and Burke [40]) for five popular DFAs. XCVerifier was successfully able to verify or find a counterexample for 11 out of the 29 (valid) DFA/condition pairs, and it was able to verify all but a small portion of the input domain for an additional 7 pairs. However, it timed out for 11 pairs, which included all the conditions for the SCAN functional. We also found that the results of the PB approach, which used grid-search to check DFT exact conditions, were inconsistent with those of XCVerifier for 5 DFA/condition pairs. These results demonstrate promise and future challenges of using formal methods for DFT.

### v5

@@
-Towards Verifying Exact Conditions of Density Functional Theory Approximations
+Towards Verifying Exact Conditions for Implementations of Density Functional Approximations
 Abstract
@@
 XCVerifier was evaluated on five DFAs from the popular Libxc library and seven exact conditions used in recent work by Pederson and Burke.
-XCVerifier was able to verify or find violations for a majority of the DFA/condition pairs, demonstrating the feasibility of using formal methods to verify DFA implementations.
-However, it timed out on all conditions of the recent SCAN functional, revealing potential directions for future work.
+XCVerifier was able to verify or find violations for a majority of the DFA-condition pairs, demonstrating the feasibility of using formal methods to verify DFA implementations.
+However, it timed out on all conditions of the recent SCAN functional, revealing directions for future work.
 Index Terms:
@@
 Density functional theory (DFT) [1, 2, 3] is a widely used approximation in the computation of electronic properties of matter.
-DFT calculations have been used to predict diverse properties–from atomic binding energies and chemical reactivities to electronic conductivities and magnetic phenomena [4].
+DFT calculations have been used to predict diverse properties from atomic binding energies and chemical reactivities to electronic conductivities and magnetic phenomena [4].
 Consequently, DFT finds applications in a number of scientific and engineering fields, including solid state physics, computational chemistry, and materials science, and has been implemented in widely-used scientific software such as ABINIT [5], cp2k [6], ERKALE [7], Psi4 [8], Octopus [9], Qbox [10], and Quantum ESPRESSO [11].
 Originating in the foundational work of Hohenberg and Kohn in 1964 [12], DFT has provided a practical way to reduce the complexity of conventional electronic structure methods by identifying the electron density function $n(r)$—a real function in three-dimensional space—as the fundamental quantity from which all other properties of a physical system can be derived.
-The Kohn-Sham (KS) [13] approach is currently the basis of most DFT calculations.
-The KS DFT approach states that only the exchange-correlation (XC) energy portion of a functional needs to be approximated.
+The Kohn-Sham (KS) [13] approach is currently the basis of most DFT calculations, and states that only the exchange-correlation (XC) energy portion of a functional needs to be approximated.
 However, the exact expression for this XC energy functional $E_{\rm xc}[n]$, which describes the complex electron-electron interactions within systems, is not known and is incredibly difficult to approximate [14].
@@
 This proliferation of DFAs is reflected in the fact that the popular Libxc software library [18], which provides numerical implementations of DFAs, currently includes over 500 functionals [19].
-The DFAs are also of varying degrees of complexity [20]: (i) local density approximations (LDAs) depend only on the electron density $n$, (ii) generalized-gradient approximations (GGAs) [21] depend on the electron density $n$ and its gradient $\nabla n$, and (iii) meta-GGAs have a further dependence on the Laplacian $\nabla^{2}n$ and the local kinetic energy density $\tau$.
+DFAs are of varying degrees of complexity [20]: (i) local density approximations (LDAs) depend only on the electron density $n$, (ii) generalized-gradient approximations (GGAs) [21] depend on the electron density $n$ and its gradient $\nabla n$, and (iii) meta-GGAs have a further dependence on the Laplacian $\nabla^{2}n$ and the local kinetic energy density $\tau$.
 The DFA $\tilde{\epsilon}_{\text{xc}}$ in Equation 1 is a mathematical function with a known, albeit complicated, analytical form.
 For instance, the following equation shows just the exchange (X) part $\tilde{\epsilon}_{\text{x}}^{\,\text{PBE}}$ of the Perdew-Burke-Ernzerhof (PBE) functional [22], a commonly used GGA functional [23]:
-$\tilde{\epsilon}_{\text{x}}^{\,\text{PBE}}(\rho,\sigma)=\frac{2.884\rho^{1/3}\left(-28.944\pi^{4/3}\rho^{8/3}-0.174\pi^{2}\sigma\right)}{\pi^{1/3}\cdot\left(77.184\pi^{4/3}\rho^{8/3}+0.257\pi^{2}\sigma\right).}$
+$\tilde{\epsilon}_{\text{x}}^{\,\text{PBE}}(\rho,\sigma)=\frac{2.884\rho^{1/3}\left(-28.944\pi^{4/3}\rho^{8/3}-0.174\pi^{2}\sigma\right)}{\pi^{1/3}\cdot\left(77.184\pi^{4/3}\rho^{8/3}+0.257\pi^{2}\sigma\right)}$
 However, the correlation part of PBE is significantly more complex with over 300 operations in the Libxc implementation.
-The SCAN meta-GGA functional [24] is even more complex with over 1000 operations in the exchange and correlation parts combined, including transcendental functions such as $\exp$ and $\log$.
-Creating a new DFA is an art that has only been mastered by few researchers as of today.
-The functional forms used in the definition of a DFA vary considerably.
+The SCAN meta-GGA functional [24] is even more complex with over 1000 operations, including transcendental functions such as $\exp$ and $\log$.
+Creating a new DFA is an art mastered by only a few researchers as of today, and the functional forms used to define the DFAs vary considerably.
 DFA designs fall into two categories: empirical and non-empirical.
@@
 For example, the correlation energy $E_{C}$ non-positivity condition states that the correlation energy cannot be positive, i.e., $E_{\rm c}[n]\leq 0$ [28].
-Furthermore, so-called norms are imposed on DFAs by requiring that they reproduce correctly some known physical systems, e.g.
-a hydrogen or a helium atom for which exact results are available.
+Furthermore, so-called norms are imposed on DFAs by requiring that they reproduce correctly some known physical systems, e.g., a hydrogen or a helium atom for which exact results are available.
 The SCAN functional is built to satisfy as many as 17 constraints and norms [24].
@@
 In particular, they checked seven exact conditions by considering their corresponding local condition (Section II).
-These local conditions were assessed for a given DFA by performing a grid search over the input variables, and checking whether each input-output pair satisfies the local condition.
+These local conditions were assessed for a given DFA by performing a grid search over the inputs to the DFA and checking whether each input-output pair satisfies the local condition.
+Many of the conditions require gradient calculations, which were numerically approximated.
 The PB approach is the state of the art in the DFT community, and was the first to perform a large scale study of the role of exact conditions in density functional development.
 For instance, they found that many empirical DFAs satisfy these exact conditions in certain regions even though they were designed without explicit adherence to these exact conditions.
-Many of the conditions require gradient or limit calculations, which need to be numerically approximated.
 This paper addresses the problem of automatically verifying whether the implementation of a DFA satisfies the exact conditions of the density functional.
 It is the first to apply formal-methods techniques to density functional theory.
-The aim is to provide formal guarantees related to the correctness of existing DFAs; viz., to formally verify if the DFA satisfies the exact conditions, and to determine the input regions in which it does not.
-The prior PB approach might miss violations of the property (because of the use of sampling) or report violations are not true (because many local conditions require gradient or limit calculations that were numerically approximated).
-As a step towards solving this problem, we designed and implemented XCVerifier, a tool that automatically converts a Libxc functional (implemented in Maple[29]) to Python, uses dReal [30] and SymPy [31] to compute any required derivatives and limits in the local condition, and then uses the dReal solver to verify that the DFA satisfies the condition in the relevant input region (Section III).
+The aim is to provide formal guarantees related to the correctness of existing DFA implementations; viz., to formally verify if the DFA implementation satisfies the exact conditions, and to determine the areas of the input domain where it does not.
+As a step towards solving this problem, we designed and implemented XCVerifier, a tool that verifies whether a Libxc functional, implemented in Maple[29], satisfies the given local condition (and, hence, the corresponding exact condition) using the dReal solver [30] (Section III).
+XCVerifier also computes any required derivatives symbolically, avoiding any potential issues arising from their numerical approximation.
 We also implemented a domain-splitting technique to improve the performance of XCVerifier (Algorithm 1).
 We evaluated XCVerifier by verifying seven exact conditions (from Pederson and Burke) for five popular DFAs: PBE [22], SCAN [24], LYP [25], AM05 [31], and VWN RPA [32], which cover the different types of DFAs (LDA, GGA, and meta-GGA), as well as different design categories (empirical and non-empirical).
-Some conditions do not apply to certain DFAs, which left us with 29 DFA/condition pairs.
-As shown in Section IV-B, XCVerifier was successfully able to verify or find a counterexample for 11 pairs, and it was able to verify all but a small portion of the input domain for an additional 7 pairs (Table I).
+Some conditions do not apply to certain DFAs, which left us with 31 DFA-condition pairs.
+As shown in Section IV-B, XCVerifier was successfully able to verify or find counterexamples for 13 pairs, and it is able to partially verify an additional seven pairs (Table I).
 These results demonstrate the feasibility of using formal-methods techniques to verify DFA implementations.
-However, XCVerifier timed out for 11 pairs: it was unable to verify the $U_{\text{c}}$ monotonicity for all but the simplest of the DFAs we considered (VWN RPA), and was unable to verify any of the properties for the SCAN DFA.
-This motivates further research on improving formal-methods techniques in the context of DFT, an important scientific computation discipline.
-We also compared the results from the PB approach with XCVerifier (Section IV-C).
-We found that the results of the PB approach were inconsistent with those of XCVerifier for 5 DFA/condition pairs, all of which were related to the two $T_{\text{c}}$ upper bound conditions (Table II).
-For instance, the PB approach found no violations to the $T_{\text{c}}$ upper bound condition for the PBE functional (Figure 2(b)), while XCVerifier was able to find a number of violations (Figure 2(e)).
-The $T_{\text{c}}$ upper bound condition entails computing derivatives and limits for which the PB approach uses numerical approximations.
-In contrast, XCVerifier computes these symbolically.
-This demonstrates one advantage of our approach to verifying DFT exact conditions.
+However, XCVerifier times out for 11 pairs: one property for the PBE DFA, three properties for AM05, and all of the properties for SCAN.
+This motivates further research on formal methods for DFT.
+To further validate our approach, we compared the results from the PB approach with XCVerifier (Section IV-C): both approaches find similar regions where the conditions were violated or satisfied (Table II).
 The contributions of the paper are as follows:
@@
   • An evaluation of XCVerifier using five DFAs and seven exact conditions along with a comparison with the state-of-the-art grid-search approach (Section IV).
-Our preliminary results demonstrate the feasibility of using formal methods to improve the correctness of DFT implementations, and reveal potential avenues for future work (Section VI).
+Our preliminary results demonstrate the feasibility of using formal methods to prove the correctness DFT implementations, and reveal avenues for future work (Section VI).
 II Exact Conditions in DFT
@@
 Each exact condition has a corresponding local condition such that if the DFA $\tilde{\epsilon}_{\text{xc}}$ satisfies the local condition, then the functional $\widetilde{E}_{\rm xc}$ satisfies the (global) exact condition.
-While the exact conditions take as input the global energy $E_{\rm xc}[n(r)]$ of the DFA, the local conditions take the exchange (correlation) enhancement factor $\widetilde{F}_{\text{xc}}$, which is a function of the local value ${\epsilon}_{\text{xc}}\big{(}n(r)\big{)}$ of the exchange (correlation) energy of the DFA.
-For GGA functionals, the inputs $n(r)$ and $\nabla n(r)$ are usually expressed in terms of the Wigner-Seitz radius $r_{s}=(r\pi n/3)^{-1/3}$ and $s=|\nabla n|/(2(3\pi^{2})^{1/3}n^{4/3})$.
-Given the expressions for the exchange and correlation energies $\tilde{\epsilon}_{\text{x}}$ and $\tilde{\epsilon}_{\text{c}}$ of the DFA in terms of $s$ and $r_{s}$, we can compute $\widetilde{F}_{\text{x}}$ and $\widetilde{F}_{\text{c}}$ of $\widetilde{F}_{\text{xc}}$:
+Note that the converse is not true: violating the local condition does not imply that the exact condition is violated.
+Furthermore, the region where the local condition is violated for a DFA can depend on the particular implementation of the functional.
+In this paper, we focus on verifying the Libxc implementation of a DFA.
+The local conditions take the exchange (correlation) enhancement factor $\widetilde{F}_{\text{xc}}$, which is a function of ${\epsilon}_{\text{xc}}\big{(}n(r)\big{)}$, the local value of the exchange (correlation) energy for the DFA.
+Here, $n(r)$ is the electron density at the point $r$ representing the position of an electron.
+For GGA functionals, the inputs $n(r)$ and $\nabla n(r)$ are usually expressed in terms of the Wigner-Seitz radius $r_{s}=(4\pi n/3)^{-1/3}$ and $s=|\nabla n|/(2(3\pi^{2})^{1/3}n^{4/3})$.
+Given the expressions for the DFA exchange and correlation energies $\tilde{\epsilon}_{\text{x}}$ and $\tilde{\epsilon}_{\text{c}}$ in terms of $s$ and $r_{s}$, we can compute $\widetilde{F}_{\text{x}}$ and $\widetilde{F}_{\text{c}}$ of $\widetilde{F}_{\text{xc}}$:
 $\widetilde{F}_{\text{xc}}[n(r)]=\widetilde{F}_{\text{x}}+\widetilde{F}_{\text{c}}=\frac{\tilde{\epsilon}_{\text{xc}}[n(r)]}{{\epsilon}_{\text{x}}^{\,\text{unif}}[n(r)]}.$ (2)
 We now list the DFT exact conditions along with their corresponding local condition:
-  • The correlation energy $E_{\rm c}$ non-positivity condition [28] is defined as
-    $E_{\rm c}[n]\leq 0.$ (3)
+(EC1) The correlation energy ($E_{\rm c}$) non-positivity condition [28] is defined as $E_{\rm c}[n]\leq 0$. The corresponding local condition is
+${\epsilon}_{\text{c}}(n(r))\leq 0.$ (3)
+It can also be expressed in terms of $\widetilde{F}_{\text{c}}$ as
+$\widetilde{F}_{\text{c}}\geq 0.$ (4)
+(EC2) The $E_{\rm c}$ scaling inequality [33, 28], $(\gamma-1)E_{\rm c}[n_{\gamma}]\geq\gamma(\gamma-1)E_{\rm c}[n]$, has the local condition
+$\frac{\partial\widetilde{F}_{\text{c}}}{\partial r_{s}}\geq 0.$ (5)
+(EC3) The $U_{\text{c}}(\lambda)$ monotonicity condition [34, 28] is $\frac{dU_{\text{c}}(\lambda)}{d\lambda}\leq 0$, where $U_{\text{c}}(\lambda)=d(\lambda^{2}E_{\rm c}[n_{1/\lambda}])/d\lambda$ represents the correlation energy adiabatic connection curves.
 The corresponding local condition is
-    ${\epsilon}_{\text{c}}(n(r))\leq 0.$ (4)
-  • The $E_{\rm c}$ scaling inequality [34, 28],
-    $(\gamma-1)E_{\rm c}[n_{\gamma}]\geq\gamma(\gamma-1)E_{\rm c}[n],$ (5)
-    has the local condition
-    $\frac{\partial\widetilde{F}_{\text{c}}}{\partial r_{s}}\geq 0.$ (6)
-  • The $U_{\text{c}}(\lambda)$ monotonicity condition [35, 28] is
-    $\frac{dU_{\text{c}}(\lambda)}{d\lambda}\leq 0,$ (7)
-    where $U_{\text{c}}(\lambda)=d(\lambda^{2}E_{\rm c}[n_{1/\lambda}])/d\lambda$ represents the correlation energy adiabatic connection curves. The corresponding local condition is
 $\frac{\partial^{2}\widetilde{F}_{\text{c}}}{\partial r_{s}^{2}}\geq\frac{-2}{r_{s}}\cdot\frac{\partial\widetilde{F}_{\text{c}}}{\partial r_{s}}.$ (6)
-  • Lieb-Oxford bound [36, 28] is
-    $U_{\text{xc}}\geq C_{\text{LO}}\int d^{3}r\,n(r)\,{\epsilon}_{\text{x}}^{\,\text{unif}}\big{(}n(r)\big{)}.$ (9)
-    where $U_{\text{xc}}[n]=E_{\rm xc}[n]-T_{\text{c}}[n]$ is the potential correlation energy and $C_{\text{LO}}=2.27$ is the Lieb-Oxford constant, following [28].
+(EC4) Lieb-Oxford bound [35, 28] is $U_{\text{xc}}\geq C_{\text{LO}}\int d^{3}r\,n(r)\,{\epsilon}_{\text{x}}^{\,\text{unif}}\big{(}n(r)\big{)}$, where $U_{\text{xc}}[n]=E_{\rm xc}[n]-T_{\text{c}}[n]$ is the potential correlation energy and $C_{\text{LO}}=2.27$ is the Lieb-Oxford constant, following [28].
 The corresponding local condition is
 $\widetilde{F}_{\text{xc}}+r_{s}\frac{\partial\widetilde{F}_{\text{c}}}{\partial r_{s}}\leq C_{\text{LO}}.$ (7)
-  • The Lieb-Oxford extension to $E_{\rm xc}$ [36, 28] is a generalization of the Lieb-Oxford bound with $E_{\rm xc}$ instead of $U_{XC}$:
-    $E_{\rm xc}\geq C_{\text{LO}}\int d^{3}r\,n(r)\,{\epsilon}_{\text{x}}^{\,\text{unif}}\big{(}n(r)\big{)}.$ (11)
+(EC5) The Lieb-Oxford extension to $E_{\rm xc}$ [35, 28] is a generalization of the Lieb-Oxford bound with $E_{\rm xc}$ instead of $U_{\text{xc}}$ :  $E_{\rm xc}\geq C_{\text{LO}}\int d^{3}r\,n(r)\,{\epsilon}_{\text{x}}^{\,\text{unif}}\big{(}n(r)\big{)}$.
 The corresponding local condition is
 $\widetilde{F}_{\text{xc}}\leq C_{\text{LO}}$ (8)
-  • The $T_{\text{c}}$ upper bound condition [35, 28] is
-    $T_{\text{c}}[n_{\gamma}]\leq-\gamma\left(\frac{\partial E_{\rm c}[n_{\gamma}]}{\partial\gamma}\right)+E_{\rm c}[n_{\gamma}].$ (13)
-    The corresponding local condition is
-    $\frac{\partial\widetilde{F}_{\text{c}}}{\partial r_{s}}\leq\frac{\widetilde{F}_{\text{c}}(\infty)-\widetilde{F}_{\text{c}}}{r_{s}}.$ (14)
-    where $\widetilde{F}_{\text{c}}(\infty)$ is the the limit of $\widetilde{F}_{\text{c}}$ as $r_{s}\to\infty$.
-  • The conjectured $T_{\text{c}}$ upper bound [37, 38, 39, 28] is
-    $T_{\text{c}}[n]\leq-E_{\rm c}[n]$ (15)
-    with local condition
+(EC6) The $T_{\text{c}}$ upper bound condition [34, 28] is $T_{\text{c}}[n_{\gamma}]\leq-\gamma\left(\frac{\partial E_{\rm c}[n_{\gamma}]}{\partial\gamma}\right)+E_{\rm c}[n_{\gamma}]$, with corresponding local condition
+$\frac{\partial\widetilde{F}_{\text{c}}}{\partial r_{s}}\leq\frac{\widetilde{F}_{\text{c}}(\infty)-\widetilde{F}_{\text{c}}}{r_{s}}$ (9)
+where $\widetilde{F}_{\text{c}}(\infty)$ is the limit of $\widetilde{F}_{\text{c}}$ as $r_{s}\to\infty$.
+(EC7) The conjectured $T_{\text{c}}$ upper bound [36, 37, 38, 28] is $T_{\text{c}}[n]\leq-E_{\rm c}[n]$ with local condition
 $\frac{\partial\widetilde{F}_{\text{c}}}{\partial r_{s}}\leq\frac{\widetilde{F}_{\text{c}}}{r_{s}}.$ (10)
 III XCVerifier
-This section describes the design the XCVerifier tool that verifies whether an DFA implemented in the Libxc library [18] satisfies the DFT exact conditions.
-XCVerifier consists of two main parts: XCEncoder, which encodes the given exact condition for a given Libxc functional into a formula $\psi$, and Verifier, which verifies whether this encoded formula $\psi$ is always true (valid) in the given input domain.
-Because DFAs involve non-linear arithmetic as well as transcendental functions, we chose the dReal solver [30] as the core solver in the Verifier.
+This section describes the design of the XCVerifier tool, which verifies whether a DFA implemented in the Libxc library [18] satisfies the DFT local conditions in Section II.
+XCVerifier consists of (i) XCEncoder, which encodes the given local condition for a given Libxc functional into a formula $\psi$, and (ii) Verifier, which verifies whether this encoded formula $\psi$ is always true (valid) in the given input domain.
+Because DFAs involve non-linear arithmetic and transcendental functions, we chose the dReal solver [30] as the core solver in Verifier.
 As a consequence, the formula $\psi$ generated by XCEncoder is a dReal formula.
 III-A XCEncoder
-Given a Libxc functional, XCEncoder first translates the Maple code for $\tilde{\epsilon}_{\text{xc}}$ to Python using CodeGeneration package of Maple.
-To test that this translation is correct, we uniformly sample points from the input region and assert that the outputs of the translated Python implementation and the original Libxc implementation are close by a factor of at least $10^{-13}$.
-Of the DFAs we use in our experiments, all points for PBE, AM05, and VWN RPA are within $10^{-14}$ of the Libxc implementation.
-LYP is within $10^{-13}$ and SCAN is within $10^{-14}$ for $98\%$ of the points sampled.
-The next step in the XCEncoder is to convert the Python code into a dReal expression.
-We implemented a symbolic execution engine for (a subset of) Python that generates this dReal expression.
-Note that though DFA implementations do not contain loops, arrays, etc., they do contain (non-recursive) function calls and if-then-else statements.
-XCEncoder needs to return the dReal formula $\psi$ that encodes the given exact condition for the particular functional.
+Given a Libxc functional, XCEncoder first translates the Maple code for $\tilde{\epsilon}_{\text{xc}}$ to Python using the CodeGeneration package from Maple.
+We implemented a symbolic execution engine for (a subset of) Python that generates the dReal expression corresponding to the DFA.
+Though DFA implementations do not contain loops, arrays, etc., they do contain (non-recursive) function calls and if-then-else statements.
+XCEncoder then constructs the dReal formula $\psi$ that encodes the given local condition for the particular functional.
 Encoding the local condition corresponding to the $E_{\rm c}$ non-positivity condition is straightforward: XCEncoder uses the dReal expression for $\tilde{\epsilon}_{\text{c}}$ to directly construct the dReal formula $\tilde{\epsilon}_{\text{c}}\leq 0$.
-However, the local conditions corresponding to exact conditions such as $E_{\rm c}$ scaling, Lieb-Oxford, $U_{\text{c}}$ monotonicity, and $T_{\text{c}}$ upper bound require computation of one or more derivatives, In such cases, XCEncoder uses the symbolic differentiation built into dReal to compute the derivatives.
-Furthermore, the local condition corresponding the $T_{\text{c}}$ upper bound condition requires computing a limit.
-In such cases, XCEncoder uses SymPy [31] to compute the limit, then translates this limit back to a dReal expression.
-After any such derivatives and limits are symbolically computed, XCEncoder constructs the required formula $\psi$ corresponding to the specific local condition.
+However, the local conditions corresponding to exact conditions such as $E_{\rm c}$ scaling, Lieb-Oxford, $U_{\text{c}}$ monotonicity, and $T_{\text{c}}$ upper bound require computation of one or more derivatives, In such cases, XCEncoder uses SymPy [39] to symbolically compute the derivatives.
+Furthermore, the local condition corresponding to the $T_{\text{c}}$ upper bound condition (Equation 9) requires computing $\lim_{r_{s}\to\infty}\widetilde{F}_{\text{c}}$.
+Following [28], XCEncoder substitutes an appropriately large value to approximate this limit at infinity, viz.
+$\widetilde{F}_{\text{c}}|_{r_{s}=100}$.
 III-B Verifier
@@
 16:  return
-The Verifier takes as input the formula $\psi$ encoding the exact condition for the DFA and a domain for the inputs to the DFA.
-We use the same input bounds as used in Pederson and Burke [40]; for example, for GGA functionals, the domain for $r_{s}$ is the interval $[0.0001,5]$, and that for $s$ is $[10^{-12},5]$ (to avoid division-by-zero errors).
-Consequently, for a GGA functional, the Verifier is trying to prove the validity of the following formula:
-$\forall r_{s},s\,.\ (r_{s}\in[0.0001,5]\wedge s\in[10^{-12},5])\implies\psi.$ (17)
-The validity of Equation 17 is equivalent to proving that the following formula is unsatisfiable:
-$r_{s}\in[0.0001,5]\wedge s\in[10^{-12},5]\wedge\lnot\psi.$ (18)
+Algorithm 1 presents the pseudo-code for Verifier.
+The Verifier component of XCVerifier takes as input (i) the formula $\psi$ encoding the exact condition for the DFA, and (ii) a domain for the inputs to the DFA.
+We use the same input bounds as used in Pederson and Burke [28]; for example, for GGA functionals, the domain for $r_{s}$ is the interval $[0.0001,5]$, and that for $s$ is $[0,5]$.
+Consequently, for a GGA functional, the Verifier is trying to prove the validity, or satisfiability, of the following formula:
+$\forall r_{s},s\,.\ (r_{s}\in[0.0001,5]\wedge s\in[0,5])\implies\psi.$ (11)
+Proving the validity of Equation 11 is equivalent to proving that the following formula is unsatisfiable:
+$r_{s}\in[0.0001,5]\wedge s\in[0,5]\wedge\lnot\psi.$ (12)
+Verifier uses the dReal solver [30] to check the satisfiability of the above formula.
 […]
-Our preliminary results showed that dReal would timeout (in 2 hours) when trying to solve the formula in Equation 18.
-To improve performance of Verifier, we implemented a domain-splitting technique that partitions the input domain and use dReal to solve the formula $\lnot\psi$ on each subdomain separately.
-This simple strategy greatly improved the performance of Verifier.
-Furthermore, one of the goals of XCVerifier is to determine the input regions where the DFA does not satisfy the exact condition.
-Thus, Verifier also performs the domain-split when dReal returns a valid input violating the given condition.
-This allows Verifier to isolate the regions where the DFA does not satisfy the local condition.
-Algorithm 1 presents the pseudo-code for Verifier, and Figure 1(e) shows a graphical representation of output of Verifier when verifying the $E_{\rm c}$ scaling inequality condition for the LYP functional.
-The dReal solver is called on Line 3 to find a satisfying assignment to the formula $\varphi_{D}\wedge\psi$, where $\varphi_{D}$ is the formula encoding the domain constraints on the inputs to the DFA.
-If the result is UNSAT, then Verifier returns after recording that the condition was verified over domain $D$ (Line 6), indicated using   in Figure 1(e).
-If the result is SAT and the returned model is a valid model for $\lnot\psi$, then Verifier records that a valid counterexample was found (Line 9), else the result is inconclusive (Line 11), indicated using   and  , respectively, in Figure 1(e).
-It is also possible that dReal is unable to determine satisfiability within the two hour time limit.
-In this case, Verifier records that the verification timed out for domain $D$ (Line 13), indicated using   in Figure 1(e).
-Line 14 is executed if the result is SAT or the solver timed out.
+Though the $\delta$-satisfiability problem is decidable in principle, in practice, the dReal solver could also timeout: it was unable to determine whether formula $\varphi$ is UNSAT or $\delta$-SAT in the given time limit.
+Our preliminary results showed that dReal would timeout on the formula in Equation 12 for most functional/condition pairs, even when given a time limit of 24 hours.
+To improve the performance of Verifier, we implement a domain-splitting technique that partitions the input domain and uses dReal to solve the formula $\lnot\psi$ on each subdomain separately.
+This simple strategy greatly improves the performance of Verifier.
+Furthermore, one of the goals of XCVerifier is to determine the input regions where the DFA implementation violates the local condition.
+Thus, Verifier also performs the domain-split when dReal returns a valid model: an input that indeed violates the given condition.
+This allows Verifier to isolate the subregions where the DFA implementation violates the local condition.
+Algorithm 1 presents the pseudo-code for Verifier, and Figure 1(f) shows a graphical representation of the output of Verifier when verifying the conjectured $T_{\text{c}}$ upper bound condition for the PBE functional.
+The dReal solver is called on Line 3 to find a satisfying assignment to the formula $\varphi_{D}\wedge\lnot\psi$, where $\varphi_{D}$ is the formula encoding the domain constraints on the inputs to the DFA and $\psi$ encodes the local condition for the functional.
+If the result is UNSAT, then Verifier returns after recording that the condition was verified over domain $D$ (Line 6).
+This is indicated using   in Figure 1(f).
+If the result is $\delta$-SAT, dReal will also return a model $x$ for the formula.
+In Line 8, valid(x) checks if this model is a valid counterexample by plugging the values back into $\psi$, which encodes the local condition.
+If the condition is indeed violated, then Verifier records it as a counterexample (Line 9), indicated using  .
+Occasionally, dReal may return SAT with a model that does not actually violate $\psi$.
+This happens due to the $\delta$-satisfiability procedure of dReal: the model must satisfy the weakened formula ($(\lnot\psi)^{\delta}$, but not necessarily the original formula $\lnot\psi$.
+In such cases, Verifier records the result as inconclusive (Line 11), indicated using   in Figure 1(f).
+We use a two hour time limit for the dReal solver; if dReal is unable to determine (un)satisfiability in this limit, Verifier interrupts it and records that the verification timed out for domain $D$ (Line 13), indicated using   in Figure 1(f).
+Increasing the timeout in our experiments did not enable dReal to solve more formulas.
+Line 14 is executed if the result is SAT or when dReal times out.
 Verifier calls split($D$), which partitions each input dimension of $D$ into two equal parts.
-The Verifier is recursively called on each subdomain $D^{\prime}$ on Line 15.
+Verifier is recursively called on each subdomain $D^{\prime}$ in Line 15.
 We set a lower limit on the size of input domain as the base case for the recursion.
-On Line 2, Verifier returns if the given input domain is too small as determined by the threshold $t$.
-For our experiments, we set the limit for recursing after a counterexample to $0.2$ for the widths of the $r_{s}$ and $s$ ranges.
-For an inconclusive or timeout result, we set it to $0.625$.
+On Line 2, Verifier returns if the given input domain is too small as determined by the threshold $t$; we used $t=0.05$ in our experiments.
 TABLE I: Verifying local conditions for DFT exact conditions for DFAs using XCVerifier.
@@
 […]
-: XCVerifier found a counterexample showing that the DFA does not satisfy the condition.
+--: the condition does not apply to the DFA;
+: XCVerifier found a counterexample showing that the DFA does not satisfy the local condition.
 Local condition PBE LYP AM05 SCAN VWN RPA
@@
 IV Experimental Results
-This section evaluates the performance of XCVerifier for verifying exact conditions for DFAs, and compares the results to the prior PB approach [28].
+This section evaluates the performance of XCVerifier for verifying local conditions for DFA implementations in Libxc, and compares the results to the prior PB approach [28].
 Our experiments were designed to answer the following research questions:
   1. RQ1
-    Is XCVerifier able to verify or find counterexamples for DFAs (Section IV-B )?
+    Is XCVerifier able to verify or find counterexamples for local conditions of DFA implementations (Section IV-B)?
   2. RQ2
-    How does compare XCVerifier to the PB approach (Section IV-C)?
+    How does XCVerifier compare to the PB approach (Section IV-C)?
 IV-A Experimental Setup
-We use the following five DFAs: PBE [22], a popular non-empirical GGA DFA; SCAN [24], a fully constrained non-empirical meta-GGA DFA satisfying all known properties of DFAs; LYP [25], and empirical DFA that is a key component of several commonly-used DFAs; AM05 [32], which shows efficient and superior performance on solids [41]; and VWN RPA [33], an LDA functional.
+We use the following five DFAs in our experiments: PBE [22], a popular non-empirical GGA DFA; SCAN [24], a fully constrained non-empirical meta-GGA DFA satisfying all known properties of DFAs; LYP [25], and empirical DFA that is a key component of several commonly-used DFAs; AM05 [31], which shows efficient and superior performance on solids [40]; and VWN RPA [32], an LDA functional.
 For each of these DFAs, we consider each of the applicable conditions from Section II.
-Note that the Lieb-Oxford conditions only apply to functionals with both an exchange and correlation component available (e.g.
-PBE, SCAN).
+Note that the Lieb-Oxford conditions only apply to functionals with both an exchange and correlation component available (e.g., PBE, AM05 SCAN).
 The PB approach is the state of the art for assessing condition satisfaction in DFT.
@@
 Table I summarizes the result of using XCVerifier to verify the local condition corresponding to each DFA exact condition in Section II for the five DFAs.
-Visualization of the results for PBE and LYP are shown in Figure 2 and Figure 1.
-From Table I, we can see that XCVerifier rarely determines that a condition is fully satisfied () for a DFA: it determines that AM05 and VWN RPA satisfy the $E_{\rm c}$ non-positivity condition, and that VWN RPA satisfies the $E_{\rm c}$ scaling inequality condition.
-For some of the DFAs, XCVerifier finds counterexamples to the conditions ().
-It finds counterexamples to the $T_{\text{c}}$ upper bound condition for all the DFAs except for SCAN, where it times out.
-It also finds counterexamples for the conjectured $T_{\text{c}}$ upper bound condition for PBE and LYP, and for all conditions but $U_{\text{c}}$ monotonicity for LYP.
-[…]
-Interestingly, XCVerifier also times out completely for every single condition applied to SCAN, which was designed to satisfy all known properties of DFAs.
-PBE  For the PBE functional, XCVerifier finds a counterexample region at $r_{s}<3.125$ and $s>1.25$ for the conjectured $T_{\text{c}}$ upper bound (Figure 2(f)).
-It is able to verify the entire input domain for the Lieb-Oxford bounds except for at $r_{s}<0.625$, where it times out.
-This is also true for the $E_{\rm c}$ non-positivity condition, which also results in some inconclusive and timeout regions.
-For the $E_{\rm c}$ scaling inequality, XCVerifier times out in the entire domain except for where $r_{s}>0.625$ and $s>3.75$.
-For the $T_{\text{c}}$ upper bound condition, it finds counterexamples in $r_{s}>0.6250875$ and times out in the rest of the region.
-LYP  For the $E_{\rm c}$ non-positivity condition, XCVerifier returns $s>1.25$ as the counterexample region (Figure 1(d)).
-For the $E_{\rm c}$ scaling inequality, it returns $r_{s}<2.19$ and $s>0.94$ (Figure 1(e)).
-It times out for the entire input domain for the $U_{\text{c}}$ monotonicity condition and for the $T_{\text{c}}$ upper bound condition, it verifies the part of the region where $1.8750.3125<r_{s}<$ 1.875 and $s>1.25$ and finds counterexamples in the rest of the region.
-For the conjectured $T_{\text{c}}$ upper bound condition, XCVerifier returns $r_{s}>0.3125$ and $s>0.9375$ as the failing region (Figure 1(f)).
-There are also some inconclusive and timeout regions for each of the conditions, usually bordering the counterexample regions.
-AM05XCVerifier verifies that AM05 satisfies $E_{\rm c}$ non-positivity for the entire region and times out completely for $U_{\text{c}}$ monotonicity.
-For the $E_{\rm c}$ scaling inequality, it verifies the entire input region except for $r_{s}<0.625$, where it times out.
-It verifies the same region for the $T_{\text{c}}$ upper bound conjectured condition and returns inconclusive for $r_{s}<0.625$.
-It finds no counterexamples for any of the conditions.
-SCAN  For the SCAN functional, XCVerifier times out after 2 hours for all of the conditions.
-To check if XCVerifier is able to verify a weaker condition for SCAN, we relax the the $E_{\rm c}$ negativity condition by adding a tolerance of $0.1$ such that the verifier needs to find a model for $E_{\rm c}<0.1$ rather than $E_{\rm c}<0$.
-However, XCVerifier times out completely for the weakened condition as well.
-VWN RPAXCVerifier verifies that VWN RPA satisfies the $E_{\rm c}$ non-positivity and $E_{\rm c}$ scaling inequality for the entire input domain.
-For the $U_{\text{c}}$ monotonicity condition, it verifies the entire region except for where $r_{s}<0.001$, where it times out.
-For the $T_{\text{c}}$ upper bound, it finds counterexamples in almost the entire region except for some small subregions in the middle of the input region where it times out.
-For the conjectured $T_{\text{c}}$ upper bound, XCVerifier is inconclusive for $r_{s}<0.3125$.
-The remaining region is verified as satisfying the condition.
-Summary for RQ1:  In our evaluation of the seven exact conditions for the five DFAs, XCVerifier was able to verify or find counterexamples for 11 condition-DFA pairs and partially verify 7, as shown in Table I.
+The results described below are summarized in Table I.
+Visualizations for LYP and PBE are shown in Figures 1 and 2, respectively.
+PBE  For the $E_{\rm c}$ non-positivity condition and the Lieb-Oxford bound, XCVerifier is able to verify the entire input domain except for some timed-out and inconclusive regions along the $s$-axis (Figure 1(d)).
+The verified region is $r_{s}>0.9375$ for the $E_{\rm c}$ non-positivity condition and $r_{s}>0.0781$ for the Lieb-Oxford bound.
+The Lieb-Oxford extension to $E_{xc}$ is verified for the entire input domain (Figure 1(e)).
+For the $E_{\rm c}$ scaling inequality, XCVerifier verifies most of the bottom third of the input domain and times out for the rest.
+There are also a few small inconclusive regions for high $s$.
+For the $T_{\text{c}}$ upper bound condition, XCVerifier verifies the bottom two-thirds of the input domain and times out for the rest.
+For the conjectured $T_{\text{c}}$ upper bound, XCVerifier finds a large counterexample region covering the upper left diagonal of the input domain (Figure 1(f)).
+There is an inconclusive region along the border of the counterexample region.
+XCVerifier verifies or times out for the rest of the input domain.
+Finally, XCVerifier times out everywhere for the $U_{\text{c}}$ monotonicity condition.
+LYP  For the $E_{\rm c}$ non-positivity condition, XCVerifier finds counterexamples at $s>1.6563$ (Figure 2(d)) and the remainder of the input domain is verified to satisfy the condition.
+For the $E_{\rm c}$ scaling inequality, the counterexamples are at $r_{s}<2.5$ and $s>1.4844$ (Figure 2(e)).
+For the $T_{\text{c}}$ upper bound condition (Figure 2(f)), the counterexamples are in a small region at $r_{s}>4.8437$ and $s>2.4219$, and for the conjectured $T_{\text{c}}$ upper bound condition, the region is $r_{s}>0.625$ and $s>1.3281$.
+For the $U_{\text{c}}$ monotonicity condition, XCVerifier finds counterexamples at $s>1.4844$ and $r_{s}<1.4062$.
+The rest of the region is partially verified or timed-out.
+There are some small inconclusive or timed-out regions on the borders of the counterexample regions for each of the conditions.
+LYP is the only DFA where XCVerifier finds counterexamples to all applicable properties.
+AM05XCVerifier verifies that AM05 satisfies the $E_{\rm c}$ non-positivity and the $T_{\text{c}}$ upper bound conditions in the entire input domain.
+It also verifies most of the domain for the $E_{\rm c}$ scaling inequality and the conjectured $T_{\text{c}}$ upper bound conditions, but times out along the $s$-axis at $r_{s}<0.0781$ for the $E_{\rm c}$ scaling inequality and $r_{s}<0.1562$ for conjectured $T_{\text{c}}$ upper bound.
+For the $U_{\text{c}}$ monotonicity condition, Lieb-Oxford bound, and Lieb-Oxford extension to $E_{xc}$, XCVerifier times out for the entire input domain.
+SCAN  For the SCAN functional, XCVerifier times out for all of the conditions.
+VWN RPAXCVerifier verifies that VWN RPA satisfies the $E_{\rm c}$ non-positivity, $E_{\rm c}$ scaling inequality, and $T_{\text{c}}$ upper bound conditions for the entire input domain.
+It also verifies the $U_{\text{c}}$ monotonicity condition where it timed out for several of the other functionals.
+For the conjectured $T_{\text{c}}$ upper bound, it verifies the entire region except along the $s$-axis at $r_{s}<0.0781$, where it returns inconclusive.
+Summary for RQ1:  In our evaluation of the seven exact conditions for the five DFAs, XCVerifier was able to verify or find counterexamples for 13 condition-DFA pairs and partially verify seven, as shown in Table I.
 This demonstrates the feasibility of using formal-methods techniques like XCVerifier in verifying DFA implementations.
-However, XCVerifier also timed out for 11 of the 29 applicable pairs, particularly for the $U_{\text{c}}$ monotonicity condition and for the SCAN functional.
-Thus, there is room for improvement in formal methods techniques for DFT.
+However, XCVerifier also timed out for 11 of the 31 applicable pairs, particularly for the SCAN functional, which was designed to satisfy all known properties of DFAs.
+Thus, there is room for improvement in formal-methods techniques for DFT.
 IV-C Comparing XCVerifier to PB
@@
 (a) $E_{\rm c}$ non-positivity w. PB
-(b) $E_{\rm c}$ scaling inequality w. PB
+(b) Lieb-Oxford ext. w. PB
 (c) Conj. $T_{\text{c}}$ upper bound w. PB
 (d) $E_{\rm c}$ non-positivity w. XCVerifier
-(e) $E_{\rm c}$ scaling inequality w. XCVerifier
+(e) Lieb-Oxford ext. w. XCVerifier
 (f) Conj. $T_{\text{c}}$ upper bound w. XCVerifier
-Figure 1: Regions where the LYP functional satisfies or violates conditions according to PB (top) and XCVerifier (bottom).
-For PB: ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\mathllap{\circ}$ is a counterexample to the condition, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bullet}\mathllap{\circ}$ is a point that satisfies the condition.
+Figure 1: Regions where the PBE functional satisfies or violates conditions according to PB (top) and XCVerifier (bottom).
+For PB:   (region hatched) is a counterexample to the condition,   is a point that satisfies the condition.
 For XCVerifier:   is a region that contains a counterexample marked with ×,   is a region that is verified to satisfy the condition,   indicates a timeout, and   indicates an inconclusive result.
 (a) $E_{\rm c}$ non-positivity w. PB
-(b) $T_{\text{c}}$ upper bound w. PB
-(c) Conj. $T_{\text{c}}$ upper bound w. PB
+(b) $E_{\rm c}$ scaling inequality w. PB
+(c) $T_{\text{c}}$ upper bound w. PB
 (d) $E_{\rm c}$ non-positivity w. XCVerifier
-(e) $T_{\text{c}}$ upper bound w. XCVerifier
-(f) Conj. $T_{\text{c}}$ upper bound w. XCVerifier
-Figure 2: Regions where the PBE functional satisfies or violates conditions according to PB (top) and XCVerifier (bottom).
-For PB: ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bullet}\mathllap{\circ}$ is a counterexample to the condition, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bullet}\mathllap{\circ}$ is a point that satisfies the condition.
+(e) $E_{\rm c}$ scaling inequality w. XCVerifier
+(f) $T_{\text{c}}$ upper bound w. XCVerifier
+Figure 2: Regions where the LYP functional satisfies or violates conditions according to PB (top) and XCVerifier (bottom).
+For PB:   (region hatched) is a counterexample to the condition,   is a point that satisfies the condition.
 For XCVerifier:   is a region that contains a counterexample marked with ×,   is a region that is verified to satisfy the condition,   indicates a timeout, and   indicates an inconclusive result.
@@
 […]
-Inconsistent results, $\bigotimes$, mean that the counterexample regions returned by PB and XCVerifier differ significantly.
-This occurs for the $T_{\text{c}}$ upper bound condition for all of the functionals except for SCAN, where XCVerifier times out.
-In our experiments, the inconsistencies for this condition are dramatic: the regions found by PB and XCVerifier are sometimes completely opposed.
-For example, for the $T_{\text{c}}$ upper bound condition for PBE (Figure 2(e)), the PB finds no counterexamples, while XCVerifier finds counterexamples in the entire region except where it times out.
-The $T_{\text{c}}$ upper bound condition is the only condition that requires a limit computation, which is done numerically in PB and symbolically in XCVerifier.
-This may explain the discrepancy.
-There is also a less dramatic inconsistency for the conjectured $T_{\text{c}}$ upper bound condition for PBE Figure 2(f).
 Several of the results are not inconsistent, marked with $\bigodot^{*}$, meaning that neither method finds counterexamples for the condition.
-This is the case for when PB finds no counterexamples and concludes that the condition is satisfied for the functional, and XCVerifier either verifies the entire region or partially times out and partially verifies the region.
-In general, this tends to looks like XCVerifier verifying the majority of the region with a smaller timed out or inconclusive region on the border of the input domain.
-This is the case for the $E_{\rm c}$ non-positivity condition (e.g.
-for PBE Figure 2(d)), the $E_{\rm c}$ scaling inequality condition, and the Lieb-Oxford conditions when XCVerifier does not time out completely.
+This is the case for when PB finds no counterexamples, and XCVerifier either verifies the entire region or partially verifies and partially times out (e.g., for PBE in Figure 1(d)).
 […]
-Summary for RQ2:  Our comparison of XCVerifier to the PB approach (Table II) shows inconsistencies for 5 DFA/condition pairs, all of which were related to the two $T_{\text{c}}$ upper bound conditions.
-The $T_{\text{c}}$ upper bound condition requires computing limits of $\widetilde{F}_{\text{c}}$, which is a function of the potentially complicated correlation energy $\tilde{\epsilon}_{\text{xc}}$.
-XCVerifier computes this limit symbolically.
-In contrast, the PB approach uses numerical approximation, which may explain the discrepancy.
-This demonstrates one advantage of our approach to verifying DFT exact conditions.
+Summary for RQ2:  The results of XCVerifier are consistent with the PB approach (Table II).
+For DFA-condition pairs for which PB finds counterexamples, XCVerifier also finds counterexamples in similar regions.
+For pairs where PB does not find counterexamples, XCVerifier either verifies the input domain, or partially times out and partially verifies the domain.
 V Related Work
+V-A Analysis of Density Functional Theory Approximations
 To the best of our knowledge, ours is the first work that uses formal methods to verify correctness in the context of density functional theory, an important scientific computing application.
-Prior work has used a testing-based approach, which are more scalable but do not provide formal guarantees.
+Prior work has used a testing-based approach, which is more scalable but does not provide formal guarantees.
 As discussed earlier, Pederson and Burke [28] use a grid-search to evaluate whether the DFA satisfies the DFT exact conditions.
-Lehtola and Marques [43] demonstrate the numerical instability of many commonly-used DFAs by observing slow convergence of self-consistent field (SCF) calculation with typically used numerical quadrature schemes.
-In general, correctness in scientific computing is recognized as a major challenge in HPC [44, 45], which needs formal methods that address the unique challenges in this domain.
+Lehtola and Marques [42] show that many recent DFAs are numerically ill-behaved by studying their accuracy in computing the total exchange-correlation energy.
+V-B Correctness in High Performance Computing
+Correctness in scientific computing is recognized as a major challenge in HPC [43, 44], which needs formal methods that address the unique challenges in this domain.
 Progress to date includes the verification of mathematical properties in a conjugate gradient solver, a finite difference stencil, and a mesh quality metric [45], PDE solvers [46], specific properties of CG [47] and LU decomposition [48], and the floating-point equivalence of manually and automatically differentiated code [49].
 The above are orthogonal to our goal of verifying exact conditions of density functional theory approximations.
-Testing and analysis techniques for floating-point programs have been developed in the past decade.
+Extending our approach of using dReal to verify properties of other scientific computations is an interesting future direction.
+V-C Analysis of Floating-Point Programs
+Many testing and analysis techniques for floating-point programs have been developed in the past decade.
 The first set of techniques are general approaches that aim to achieve high-coverage of numerical code  [50, 51], conduct differential testing of numerical libraries [52], and perform mutation testing of floating-point expressions given a real specification [53].
-However, it has been shown that simply achieving high code coverage in numerical programs does not uncover numerical issues in most cases [52], and differential testing is not feasible for DFAs because most of them are unique.
+However, it has been shown that simply achieving high code coverage in numerical programs does not uncover numerical issues in most cases [51].
+Additionally, differential testing is not feasible for DFA implementations because most of them are unique.
 A large body of work focuses on performing automated error analysis of floating-point programs [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65].
 While some of these approaches provide sound error bounds of floating-point programs, they suffer from important limitations with respect to program size and control structures supported.
-In the absence of techniques that can reason about floating-point error in non-trivial programs, a rich area of research in software testing has focused on how to efficiently generate inputs that maximize error in the output of a program  [67, 52, 68], which can shed light on the potentially worst error a floating-point program could incur in.
+In the absence of techniques that can reason about floating-point error in non-trivial programs, a rich area of research in software testing has focused on how to efficiently generate inputs that maximize error in the output of a program [66, 51, 67], which can shed light on the potentially worst error a floating-point program could incur.
 Similarly, work has proposed techniques to generate inputs that trigger floating-point exceptions [68, 69].
 Calculating error bounds of DFAs is orthogonal to our goal of verifying physical and numerical properties of their implementations.
-Existing work has explored finding function input ranges, also referred to as regimes, with the purpose of improving the accuracy of floating-point expressions [71] or optimizing floating-point efficiency [72].
+Other existing work on floating-point programs has explored finding function input ranges, also referred to as regimes, with the purpose of improving the accuracy of floating-point expressions [70] or optimizing floating-point efficiency [71].
 These approaches are based on either estimating error based on dynamic input sampling, or statically performing an error analysis, which have their own limitations, as described earlier.
@@
 VI-A Improving Scalability of the Solver
-XCVerifier was unable to verify any of the exact conditions of the SCAN functional, which has been designed to satisfy all known exact conditions.
+In our experiments, XCVerifier was unable to verify any of the exact conditions of the SCAN functional, which has been designed to satisfy all known exact conditions.
 SCAN is significantly more complex than the other functionals we considered, and also involves the use of transcendental functions such as $\exp$ and $\log$.
-This causes dReal to timeout even when the input domain is reduced.
+This causes dReal to time out even for the relatively simple $E_{\rm c}$ non-positivity condition, and even when the input domain is reduced 32×32$\times$32 ×.
 It would be interesting to investigate approaches to improve the performance of the solver so that it can tackle the SCAN functional.
-Apart from its popularity, the SCAN functional will serve as a fascinating use case: there are a progression of DFAs—rSCAN, r++SCAN, r2SCAN, r4SCAN—proposed in the literature that were designed with different adherence to exact conditions to improve the numerical stability of the original SCAN functional [73, 74, 75].
-VI-B Expanding to more DFA/condition pairs
+Apart from its popularity, the SCAN functional will serve as a fascinating use case: there is a progression of DFAs—rSCAN, r++SCAN, r2SCAN, r4SCAN—proposed in the literature that were designed with different adherence to exact conditions to improve the numerical stability of the original SCAN functional [72, 73, 74].
+VI-B Expanding to More DFA-Condition Pairs
 Our evaluation demonstrated the robustness of the XCEncoder to handle a wide variety of DFAs and exact conditions.
 The ultimate goal of our research is to be able to analyze all the 500+ functionals in Libxc for all known DFT exact conditions.
-Future work will continue to expand our evaluation, and will aim to integrate our verification tool into the Libxc, e.g., as part of the continuous integration (CI) for Libxc.
+Future work will continue to expand our evaluation, and will aim to integrate our verification tool into Libxc, e.g., as part of the continuous integration (CI) for Libxc.
 VI-C Numerical Issues With DFAs
-Apart from verifying known exact conditions for DFA implementations, it would interesting to analyze numerical issues of the implementations with the goal of using formal methods to find and fix numerical issues in DFA implementations.
+Apart from verifying known exact conditions for DFA implementations, it would be interesting to analyze numerical issues of the implementations with the goal of using formal methods to find and fix numerical issues in DFA implementations.
 This is a challenging problem involving reasoning about floating points and dealing with transcendental functions like $\sin$, $\log$ and $\exp$.
@@
 The functional form of a DFA may also make it sensitive to inaccuracies in its input data.
-While a given implementation of a DFA may yield correct answers for an exactly known (e.g.
-exponential) density, it may result in large numerical errors if the input density is noisy or if the density and its gradient are not numerically consistent.
-This is particularly problematic in regions of low density, e.g.
-far from a molecule placed in vacuum.
+While a given implementation of a DFA may yield correct answers for an exactly known (e.g., exponential) density, it may result in large numerical errors if the input density is noisy, or if the density and its gradient are not numerically consistent.
+This is particularly problematic in regions of low density, e.g., a point far from a molecule placed in vacuum.
 Such large errors may lead to inaccurate energies or slow convergence in the solution of the Kohn-Sham equations.
@@
 In other cases, an analytical reformulation of a DFA is used to avoid numerical issues [77] without modifying it.
-These fixes are however ad hoc, and there is no known general recipe for avoiding the numerical issues of a DFA.
+However, these fixes are ad hoc, and there is no known general recipe for avoiding the numerical issues of a DFA.
 VII Conclusion
-This paper presented XCVerifier, an approach for verifying whether a DFA implementation satisfies DFT exact conditions.
-XCVerifier automatically encodes the DFA implementation from Libxc and a given exact condition into a dReal formula, symbolically performing any required derivative and limit calculations.
+This paper presented XCVerifier, an approach for verifying whether a DFA implementation satisfies the DFT exact conditions.
+XCVerifier automatically encodes the DFA implementation from Libxc and a given exact condition into a dReal formula, symbolically performing any required derivative calculations.
 It uses the dReal solver to verify whether the condition is true or find a violation to it for a given input domain.
-XCVerifier also implements a domain-splitting technique to improve performance by reducing solver timeouts, and isolate the input regions where the condition is satisfied or violated.
+XCVerifier also implements a domain-splitting technique to improve performance, reducing solver timeouts, and isolating the input regions where the condition is satisfied or violated.
 We evaluated XCVerifier by verifying seven exact conditions (from Pederson and Burke [28]) for five popular DFAs.
-XCVerifier was successfully able to verify or find a counterexample for 11 out of the 29 (valid) DFA/condition pairs, and it was able to verify all but a small portion of the input domain for an additional 7 pairs.
+XCVerifier was successfully able to verify or find a counterexample for 13 out of the 31 (valid) DFA-condition pairs, and it was able to partially verify an additional seven pairs.
 However, it timed out for 11 pairs, which included all the conditions for the SCAN functional.
-We also found that the results of the PB approach, which used grid-search to check DFT exact conditions, were inconsistent with those of XCVerifier for 5 DFA/condition pairs.
+We found that the results of the PB approach, which used grid search to check DFT exact conditions, were largely consistent with those of XCVerifier.
 These results demonstrate promise and future challenges of using formal methods for DFT.